Speedy Auto Repairs uses a job-order costing system. The company’s direct materials consist of replacement parts installed in customer vehicles, and its direct labor consists of the mechanics’ hourly wages. Speedy’s overhead costs include various items, such as the shop manager’s salary, depreciation of equipment, utilities, insurance, and magazine subscriptions and refreshments for the waiting room.

The company applies all of its overhead costs to jobs based on direct labor-hours. At the beginning of the year, it made the following estimates:

Direct labor-hours required to support estimated output 38,000
Fixed overhead cost $ 589,000
Variable overhead cost per direct labor-hour $ 1.00

Required:

1. Compute the predetermined overhead rate.

2. During the year, Mr. Wilkes brought in his vehicle to replace his brakes, spark plugs, and tires. The following information was available with respect to his job:

Direct materials $ 617
Direct labor cost $ 221
Direct labor-hours used 8

Compute Mr. Wilkes’ total job cost.

3. If Speedy establishes its selling prices using a markup percentage of 40% of its total job cost, then how much would it have charged Mr. Wilkes?

Compute the predetermined overhead rate. (Round your answer to 2 decimal places.)

Predetermined overhead rate per DLH

Compute Mr. Wilkes’ total job cost. (Round your intermediate calculations to 2 decimal places.)

Direct materials
Direct labor
Overhead applied
Total cost assigned to Mr. Wilkes

If Speedy establishes its selling prices using a markup percentage of 40% of its total job cost, then how much would it have charged Mr. Wilkes? (Round your intermediate calculations to 2 decimal places.)

Amount charged to Mr.Wikes
Answer #1

1) Predetermine overhead rate = 589000/38000+1 = 16.50 per labor hour

2) Total Job Cost

Direct material 617
Direct labor 221
Overhead (16.50*8) 132
Total Job Cost 970

3) Selling price = 970*1.4 = $1358